### v1

# Controllable Logical Hypothesis Generation for Abductive Reasoning in Knowledge Graphs

Abstract
Abductive reasoning in knowledge graphs aims to generate plausible logical hypotheses from observed entities, with broad applications in areas such as clinical diagnosis and scientific discovery. However, due to a lack of controllability, a single observation may yield numerous plausible but redundant or irrelevant hypotheses on large-scale knowledge graphs. To address this limitation, we introduce the task of controllable hypothesis generation to improve the practical utility of abductive reasoning. This task faces two key challenges when controlling for generating long and complex logical hypotheses: hypothesis space collapse and hypothesis oversensitivity. To address these challenges, we propose CtrlHGen, a Controllable logcial Hypothesis Generation framework for abductive reasoning over knowledge graphs, trained in a two-stage paradigm including supervised learning and subsequent reinforcement learning. To mitigate hypothesis space collapse, we design a dataset augmentation strategy based on sub-logical decomposition, enabling the model to learn complex logical structures by leveraging semantic patterns in simpler components. To address hypothesis oversensitivity, we incorporate smoothed semantic rewards including Dice and Overlap scores, and introduce a condition-adherence reward to guide the generation toward user-specified control constraints. Extensive experiments on three benchmark datasets demonstrate that our model not only better adheres to control conditions but also achieves superior semantic similarity performance compared to baselines.
1 Introduction
Abduction is widely recognized as one of the three major types of reasoning in philosophy abduction . Specifically, abductive reasoning abduction is a form of logical inference that seeks the best or most plausible hypothesis to explain an observed phenomenon and it plays a vital role across various fields abducreview . For example, it serves as a critical tool for hypothesizing causal links between symptoms and underlying pathologies in clinical diagnosis medicaldianosis ; medical2 . Similarly, abductive methods localize system faults by interpreting anomalous signal patterns in anomaly detection attack ; attack2 . Its power also extends to scientific discovery science1 ; science2 ; science3 ; science4 , including the deduction of unknown celestial bodies from gravitational perturbations in orbital trajectories sciencediscovery .
(a) Control Semantic Domain
(b) Control Structure Complexity
Figure 1: Examples of Controllability in Abductive Reasoning
On the other hand, effective abductive reasoning requires high-quality interconnected information as premises. Knowledge graphs, whether general or domain-specific, provide an ideal structure for these premises, supporting robust abductive reasoning processes. Abductive reasoning in knowledge graphs aims to generate complex logical hypotheses that best explain a given set of observed entities, while maintaining consistency with the existing knowledge base. This process leverages the rich domain knowledge, thereby improving both the precision and reliability of the inference. AbductiveKGR akgr is the first work to realize abductive reasoning over knowledge graphs, training a hypothesis generation model using a supervised learning–reinforcement learning paradigm.
However, as knowledge graphs often contain millions of facts, an observation can lead to numerous diverse yet plausible hypotheses, most of which may be irrelevant to the specific concerns. To address this challenge, we are motivated to introduce controlling mechanisms in the hypothesis generation process that focus on two critical aspects:
Controlling semantic content enables aspect-specific reasoning. We prioritize semantic control to narrow vast hypothesis spaces to relevant aspects, essential for specialized fields where aspect-specific insights drive decision-making. As shown in Fig. 1(a), we want to explain the observation involving three diseases:{Systemic Lupus Erythematosus, Antiphospholipid Syndrome, and Sjögren’s Syndrome}. Directing attention to specific aspects—such as pathology, treatment, or affected populations—yields hypotheses that are precisely aligned with each aspect. From the pathology aspect, these diseases produce autoantibodies and are both chronic and autoimmune. From the treatment aspect, these autoimmune diseases can be treated with hydroxychloroquine. Finally, from the susceptibility aspect, these diseases are more likely to occur in women of reproductive age who carry the HLA-DR3 susceptible allele. Although these hypotheses are all plausible, their usefulness varies when people seek explanations for different scenarios.
Controlling structural complexity adjusts the level of granularity. We focus on complexity control to address varying information needs across different reasoning scenarios and align with users’ cognitive preferences for adjustable information density. In Fig. 1(b), for an observation composed of three NBA players, increasing the complexity of the hypothesis structure enables the model to capture richer shared experiences or achievements among them. By adjusting the structural complexity, users can flexibly decide how much information they want to include in the generated hypotheses. Unfortunately, prior work akgr has largely overlooked controllable generation, resulting in hypotheses that are redundant or lack meaningful relevance.
Motivated by these, we introduce the task of controllable abductive reasoning, aiming at controllable generation of hypothesis, which leads to better leverage the practical value of abductive reasoning in knowledge graphs. However, when implementing semantic and structural controls on complex logical hypotheses, we face two critical challenges: (i) Hypothesis Space Collapse: As illustrated in Fig. 2(a), the number of plausible hypotheses drops sharply as their length increases. This steep decline severely constrains our ability to apply structural complexity controls, as the available hypothesis space for longer, more complex structures becomes extremely limited. When attempting to simultaneously satisfy semantic content controls or structural complexity requirements, the intersection of valid hypotheses may become vanishingly small. (ii) Hypothesis Oversensitivity: As shown in Figure 2(b), when generating long logical hypotheses under control conditions, even small errors can result in significant differences in the conclusions. Previous work akgr used the Jaccard index as a reward in reinforcement learning, enforcing a perfect match between the hypothesis conclusion and the observation. This strict requirement may cause the model to deviate from the control conditions in order to maintain reward stability.
To tackle these challenges, we propose a Controllable logcial Hypothesis Generation method (CtrlHGen) for abductive reasoning in knowledge graphs. To address the problem of hypothesis space collapse, we introduce a dataset augmentation strategy based on sub-logical decomposition. By leveraging the semantic similarity of simpler sub-logics derived from the decomposition of complex hypotheses, this approach enables the model to understand long logical structures, which are composed of these smaller components. The hypothesis generator is then trained using a combination of supervised fine-tuning and reinforcement learning. To address the problem of hypothesis oversensitivity, we refine the semantic reward function by incorporating the Dice coefficient to smooth out minor discrepancies between the hypothesis and the target. Additionally, we introduce a condition-adherence reward to encourage the generation of hypotheses that adhere to the control constraints during exploration. Our main contributions are as follows:
  • We are the first to introduce the task of controllable abductive reasoning, enabling abductive reasoning in knowledge graphs to better satisfy practical needs by controlling semantic content and structural complexity.
  • We propose an observation-hypothesis pair augmentation strategy via sub-logical decomposition to address the challenge of hypothesis space collapse when generating complex logical structures, significantly enhancing the quality of controllable hypotheses.
  • To mitigate hypothesis oversensitivity, we refine the semantic reward function by incorporating Dice and Overlap coefficients to accommodate minor discrepancies between hypotheses and targets, while introducing a condition-adherence reward to ensure better compliance with control constraints, leading to more stable and accurate learning.
  • Extensive experiments on three datasets demonstrate that our model not only adheres more effectively to control signals but also achieves superior semantic similarity performance compared to the baseline across multiple evaluation metrics.
(a) Hypothesis Space Collapse
(b) Hypothesis Oversensitivity
Figure 2: (a) Hypothesis quality (measured in Jaccard) and space size across three logic lengths: short (one predicate), medium (two predicates), and long (three predicates). Valid candidates represent average reference hypotheses per observation. Note the dramatic collapse of hypothesis space as complexity increases. (b) Hypothesis oversensitivity example: Minor errors cause significant Jaccard score drops, creating tension between control adherence and semantic accuracy.
2 Related Work
Knowledge Graph Reasoning. Deductive reasoning focuses on answering complex logical queries by improving query and answer embeddings multihop ; query2box ; query2particles ; bai2023complex ; bai2023knowledge ; bai2024understanding . Inductive reasoning, often framed as rule mining, ranges from efficient symbolic methods like AMIE amie to embedding-based approaches such as RuLES rules and RLogic Rlogic , though traditional search-based techniques face scalability challenges. Abductive reasoning was introduced by AbductiveKGR akgr using Transformer-based hypothesis generation, with follow-up work ngdb highlighting its future potential.
Abductive Reasoning. In natural language inference, $\alpha$-NLI bhagavatula2020abductivecommonsensereasoning introduced abductive reasoning to commonsense reasoning, where plausible explanations are inferred from observations. Subsequent works proposed various techniques to enhance this capability qin2021futureunsupervisedbackpropbaseddecoding ; kadiķis2022embarrassinglysimpleperformanceprediction ; chan2023selfconsistentnarrativepromptsabductive , including extensions to uncommon scenarios focusing on rare but logical explanations zhao2024uncommonsensereasoningabductivereasoning . Unlike real-world data in commonsense reasoning, benchmarks like ProofWriter tafjord2021proofwritergeneratingimplicationsproofs evaluate formal abductive reasoning within semi-structured texts with explicit logical relationships. Recent studies have explored LLMs in more challenging open-world reasoning contexts zhong2023chatablabductivelearningnatural ; del2023truedetectivedeepabductive ; thagard2024chatgptmakeexplanatoryinferences and abstract reasoning tasks liu2024incompleteloopinstructioninference ; zheng2025logidynamicsunravelingdynamicslogical .
Meanwhile, in the neuro-symbolic domain, Abductive Learning (ABL) ABL attempts to integrate machine learning and logical reasoning in a balanced and mutually supportive manner. Recent research in this area, exemplified by systems such as ARLC ABL2 and ABL-Refl ABL3 , focuses on enhancing this integration by introducing novel techniques to improve context-awareness, error correction, generalization, and overall reasoning accuracy and efficiency.
3 Method
Figure 3: An overview of our controllable abductive reasoning framework. The process consists of three main steps: (1) Hypothesis-Observation pair construction through sub-logic decomposition to expand the hypothesis space, (2) Supervised training of the generative model using augmented hypotheses, and (3) Reinforcement tuning with dual rewards for semantic alignment and condition adherence to balance hypothesis accuracy with control signal compliance.
In this section, we elaborate the proposed CtrlHGen, a controllable hypothesis generation method for abductive reasoning in knowledge graphs. The framework of CtrlHGen is shown in Fig. 3.
3.1 Problem Definition
We define a knowledge graph as $G=(V,R)$, where $V$ is the set of entities and $R$ is the set of binary relations. A triple $(u,r,v)$ exists in $G$ if $r(u,v)=\text{true}$. Following the open-world assumption openworld , only the observed graph $G$ is available during training, with missing triples treated as unknown rather than false. The full graph $\bar{G}$ remains hidden, and $G\subseteq\bar{G}$.
The core concepts of abductive reasoning consist of observation and hypothesis. Here, an observation $O$ in knowledge graph $G$ is defined as a set of entities $O=\{o_{1},o_{2},\ldots,o_{n}\}$, where $o_{i}\in V,\forall i\in\{1,\ldots,n\}$ A logical hypothesis $H$ is defined as a query in the form of first-order logic on a knowledge graph $G$, including existential quantifiers($\text{And}(\wedge),Or(\vee),Not(\neg)$). The hypothesis can also be written in disjunctive normal form:
$H(V_{?})$ $=\exists V_{1},\ldots,V_{k}:e_{1}\vee\cdots\vee e_{n},$ (1)
$e_{i}$ $=r_{i1}\wedge\cdots\wedge r_{im_{i}},$
where $\{V_{1},\ldots,V_{k}\}$ denotes the subset of $V$. Each $r_{ij}$ is defined as either $r_{ij}=r(u,v)$ or $r_{ij}=\neg r(u,v)$, where $u$ and $v$ are either fixed entities from the set $\{V_{1},\ldots,V_{k}\}$, or variable vertices $V_{?}$, which could be any entity on the graph $G$.
The conclusion of the hypothesis $[H]_{G}$ on a graph $G$ is the set of the variable entities $V_{?}$ for which $H$ holds true on $G$. Specifically, it can be formulated as:
$[H]_{G}=\{V_{?}\in G|H(V_{?})=\mathrm{true}\}.$ (2)
Definition 3.1 (Controllable Abductive Reasoning in Knowledge Graph).
Given a knowledge graph $G$, an observation $O$, and a control condition $C$, the goal of controllable abductive reasoning is to find a hypothesis $H$ satisfying:
  1. 1.
    The hypothesis $H$ is the most plausible explanation for the observation $O$ In other words, the conclusion $[H]_{G}$ closely matches the observation $O$.
  2. 2.
    $H$ satisfies the constraints specified by the control condition $C$.
3.2 Observation-Hypothesis Pairs Construction
Sampling. We randomly sample observation-hypothesis pairs from the knowledge graph by constructing hypotheses based on predefined logical patterns. Each logical pattern is assigned an equal number of hypotheses to ensure diversity, and the query results(conclusion) of hypotheses on the graph are taken as the corresponding observations. Finally, both hypotheses and observations are converted into input sequences suitable for the generative model.
Augmentation by sub-logic decomposition. To address the challenge of hypothesis space collapse in complex logical patterns, we propose a dataset augmentation method based on sub-logic decomposition. Specifically, given a hypothesis–observation pair $(H,O)$ under a complex logical pattern $P$, we recursively decompose the hypothesis into sub-hypotheses $H_{\text{sub}}$ according to identifiable sub-logical patterns $P_{\text{sub}}$. Corresponding sub-observations $O_{\text{sub}}$ are then derived by executing these sub-hypotheses on the knowledge graph $G$. This process effectively generates additional hypothesis–observation pairs and can be formally described as:
$\{(H_{\text{sub}}^{i},O_{\text{sub}}^{i})\}_{i=1}^{n}=\left\{(f(P_{\text{sub}}^{i},H),[f(P_{\text{sub}}^{i},H)]_{G})\;\middle|\;P_{\text{sub}}^{i}\subseteq P\right\},$ (3)
where $f(P_{\text{sub}}^{i},H)$ denotes the sub-hypothesis generated based on the sub-pattern $P_{\text{sub}}^{i}$ and the origin hypothesis $H$, and $[f(P_{\text{sub}}^{i},H)]_{G}$ computes the corresponding sub-observation by querying the knowledge graph to get the conclusion of the sub-hypothesis.
Because each sub-hypothesis is a subset of the original, they are closely related both structurally and semantically. This strong alignment enables the model to progressively learn complex logical patterns by building on simpler, related sub-patterns. We have reported more details and examples for the observation-hypothesis pairs construction process in Appendix A.
3.3 Supervised Training of Controllable Hypothesis Generation
To enable controllable generation of logical hypotheses, we train a conditional generative model to generate hypothesis sequences guided by a given observation and control condition. Specifically, given an observation sequence $O=\{o_{1},\ldots,o_{m}\}$, a target hypothesis sequence $H=\{h_{1},\ldots,h_{n}\}$, and a control condition $C$, the generative model is optimized using an autoregressive loss:
$\mathcal{L}_{\text{AR}}$ $=-\log p_{\theta}(H\mid O,C)=-\sum\nolimits_{i=1}^{n}\log p_{\theta}(h_{i}\mid h_{1},\ldots,h_{i-1},O,C),$ (4)
where $\theta$ denotes the generative model, which we implement using a standard Transformer-based decoder-only architecture.
The training process consists of two stages. In the first stage, the model is trained under an unconditional setting, where the input only consists of observation tokens. This allows the model to acquire a general capability for hypothesis generation. In the second stage, the model is fine-tuned under different control conditions respectively. The input is formed by concatenating observation tokens with control condition tokens, guiding the model to generate hypotheses that satisfy the constraints.
The control conditions $C$ are designed from two different perspectives to guide hypothesis generation:
  • Semantic Focus: We randomly sample a specific entity or relation from the target hypothesis as a control condition. This guides the model to generate hypotheses grounded in a specific semantic region of the knowledge graph. The control condition is directly represented by the token of the selected entity or relation. Formally, $C\in\{T_{e}\}$ or $C\in\{T_{r}\}$. $T_{e}$ and $T_{r}$ represents the token set of entity and relation respectively.
  • Structural Constraint: We apply constraints based on the logic structure of the hypothesis. Specifically, we implement three types of structural control: (1) strictly enforcing a predefined logical pattern, where each logical pattern is represented in Lisp-like language with operator tokens following previous work in KG reasoning baisequential ; akgr . (2) constraining the number of entities involved, encoded using a special token [ne] that indicates hypotheses with exactly $n$ entities. Formally, $C\in\{[ne]\}$, where $n$ is an Integer. (3) constraining the number of relations used in the generated hypothesis, encoded using a token [nr], where [nr] denotes hypotheses containing exactly $n$ relations. Formally, $C\in\{[nr]\}$, where $n$ is an Integer.
3.4 Reinforcement Learning
To improve the generalization ability on unseen knowledge graphs and better adhere to the specified control conditions, we further fine-tune the generative model using reinforcement learning. The reward function is constructed from two perspectives: semantic alignment and condition adherence.
Semantic Alignment: This reward assesses the semantic consistency between the generated hypothesis conclusion $[H]_{G}$ and the corresponding observation $O$. We adopt the Jaccard similarity coefficient as the primary reward due to its strict evaluation of set-level agreement. However, the high sensitivity of hypotheses can lead to sharp reward fluctuations in response to minor errors. To mitigate this, we integrate two supplementary metrics:the Dice similarity coefficient and the Overlap similarity coefficient, which provide smoother gradients and greater tolerance to slight mismatches. The final semantic reward $R_{\text{sem}}$ is computed as a weighted combination of these metrics, defined as:
$R_{\text{sem}}([H]_{G},O)$ $=\lambda_{1}\cdot\text{Jaccard}([H]_{G},O)+\lambda_{2}\cdot\text{Dice}([H]_{G},O)+\lambda_{3}\cdot\text{Overlap}([H]_{G},O)$ (5)
$=\lambda_{1}\cdot\frac{|[H]_{G}\cap O|}{|[H]_{G}\cup O|}+\lambda_{2}\cdot\frac{2|[H]_{G}\cap O|}{|[H]_{G}|+|O|}+\lambda_{3}\cdot\frac{|[H]_{G}\cap O|}{\min(|[H]_{G}|,|O|)},$
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters. $G$ denotes the observable knowledge graph during training, which serves as a reliable and leakage-free proxy for evaluating abductive reasoning quality.
Condition Adherence: This reward encourages the model to generate hypotheses that satisfy the given control condition $C$. We formulate it as a binary-valued function: if the generated hypothesis $H$ satisfies the condition $C$, the reward is 1; otherwise, it is 0. The final adherence performance is evaluated by computing the proportion of generated hypotheses that meet the condition. Formally, the reward function is defined as:
$R_{\text{cond}}(H,C)=\begin{cases}1,&\text{if }H\text{ satisfies }C,\\ 0,&\text{otherwise}.\end{cases}$ (6)
Jointly capturing condition adherence and semantic alignment, the overall reward function $\hat{R}$ is formulated as:
$\hat{R}(H,O,C,G)=\alpha\cdot R_{\text{sem}}([H]_{G},O)+(1-\alpha)\cdot R_{\text{cond}}(H,C),$ (7)
where $\alpha\in[0,1]$ is a hyperparameter that balances the contributions of semantic alignment and condition adherence.
Since abductive reasoning often involves generating multiple plausible hypotheses rather than a single answer, it is important to ensure overall hypothesis quality. To achieve this, we use Group Relative Policy Optimization (GRPO) GRPO , which promotes consistent improvement across a set of sampled hypotheses per observation, instead of optimizing individual outputs. Specifically, GRPO updates the model $\pi_{\theta}$ by maximizing the expected reward over a group of hypotheses $\hat{H}={H_{1},\ldots,H_{k}}$ sampled from the same observation $O$ and control condition $C$. The objective is:
$\mathcal{J}(\theta)=\mathbb{E}_{O,\{H_{i}\}\sim\pi_{\theta_{\text{old}}}(H|O,C)}[\frac{1}{k}\sum_{i=1}^{k}\frac{1}{|H_{i}|}\sum_{t=1}^{|H_{i}|}\left\{\frac{\pi_{\theta}(h_{i,t}|O,C,h_{i,<t})}{\pi_{\theta_{\text{old}}}(h_{i,t}|O,C,h_{i,<t})}\hat{R}^{{}^{\prime}}_{i}-\beta\mathrm{D}_{KL}\left[\pi_{\theta}||\pi_{\text{ref}}\right]\right\}],$ (8)
where $k$ is the number of sampled hypotheses per observation. The normalized reward $\hat{R}^{\prime}_{i}$ is obtained by applying intra-group normalization over $\{\hat{R}_{1},\ldots,\hat{R}_{k}\}$. A KL term constrains the policy $\pi_{\theta}$ from drifting too far from the reference model $\pi_{\text{ref}}$, with $\beta$ controlling its strength. Gradient clipping is also used to stabilize training.
4 Experiment
4.1 Experiment Settings
Dataset. We conduct experiments on three widely used knowledge graph datasets: DBpedia50 dbpedia , WN18RR wn18rr , and FB15k-237 fb15k . Following akgr , each dataset is split into training, validation, and test sets with an 8:1:1 ratio. Under the open-world assumption, we incrementally build $G_{\text{train}}$, $G_{\text{valid}}$, and $G_{\text{test}}$, where each graph includes all previous edges. Additional dataset details are provided in Appendix B.
Observation-Hypothesis Pair. Following prior KG reasoning work akgr , we adopt the 13 predefined logical patterns in Fig. 4 for hypothesis sampling. Each observation contains no more than 32 entities. To evaluate generalization, the validation and test sets include entities not seen during training, with the test set covering more unseen entities. For data augmentation, we decompose five complex logical patterns (up, 3in, pni, pin, inp). Detailed sampling statistics are provided in Appendix B.
Figure 4: Thirteen predefined logical types.
Evaluation Metrics. The quality of generated hypotheses is evaluated in terms of semantic similarity and condition adherence. For semantic similarity, we use Jaccard, Dice and Overlap index, with $G_{\text{test}}$ used to compute $[H]_{G_{\text{test}}}$ during testing. For condition adherence, we regard it as a binary classification problem and calculate Accuracy. In addition, Smatch score is also used to quantify the structural similarity corresponding to the generated hypothesis $H$ and the reference hypothesis $H_{\text{ref}}$. It can measure how similar the nodes, edges and their labels are by representing the hypothesis as a graph. It should be noted that Smatch is only a reference metric, as the generated hypotheses do not need to be the same as the reference hypotheses. More information about Smatch is reported in Appendix B.
Implementation Details. We adopt a GPT-2 gpt2 architecture for the hypothesis generation model and use the AdamW optimizer. All experiments are conducted on 4 Nvidia A6000 48GB GPUs. Additional hyperparameter settings and a discussion on why LLMs are not employed are provided in Appendix B  [^1] The code is available at https://github.com/HKUST-KnowComp/CtrlHGen ..
Table 1: The results of controllable abductive reasoning under different conditions. (Result: average score $\pm$ standard deviation. Bold: best; Underline: runner-up. —: cannot be evaluated.)
Dataset Condition Semantic Similarity Condition Adherence
Jaccard Dice Overlap Accuracy Smatch
FB15k-237 uncondition 61.4±0.33 69.3±0.31 82.3±0.33 61.4±0.21
pattern 65.5±0.33 73.0±0.30 83.9±0.27 98.9±0.10 82.3±0.10
relation-number 65.1±0.33 72.7±0.31 83.5±0.29 99.4±0.14 82.4±0.20
entity-number 63.1±0.33 71.5±0.30 82.7±0.28 86.3±0.02 65.7±0.10
specific-entity 64.3±0.35 71.1±0.33 82.4±0.31 98.9±0.10 71.2±0.21
specific-relation 63.3±0.34 71.4±0.32 82.6±0.30 99.5±0.06 64.8±0.21
WN18RR uncondition 72.6±0.35 74.2±0.33 85.2±0.31 56.4±0.20
pattern 77.0±0.34 80.8±0.31 86.8±0.28 93.5±0.24 83.3±0.15
relation-number 74.0±0.34 77.4±0.31 86.3±0.28 95.3±0.25 78.9±0.20
entity-number 73.2±0.37 77.9±0.35 87.2±0.33 85.2±0.28 65.1±0.18
specific-entity 73.6±0.38 75.6±0.37 86.2±0.36 89.0±0.31 65.2±0.21
specific-relation 73.0±0.35 75.2±0.33 85.7±0.30 96.1±0.19 60.8±0.21
DBpedia50 uncondition 64.3±0.35 66.2±0.33 79.5±0.30 51.0±0.24
pattern 73.8±0.37 76.6±0.36 86.8±0.26 88.4±0.36 79.2±0.20
relation-number 72.1±0.32 76.1±0.30 87.5±0.22 80.6±0.43 79.1±0.22
entity-number 75.2±0.37 80.3±0.35 92.4±0.29 84.0±0.26 63.3±0.22
specific-entity 73.7±0.33 78.7±0.31 88.4±0.35 79.6±0.40 62.9±0.22
specific-relation 75.2±0.31 80.6±0.29 93.7±0.20 84.2±0.36 60.3±0.20
4.2 Experiment Results and Analysis
We evaluated the quality and controllability of generated hypotheses on three datasets under five conditions: pattern, relation-number, entity-number, specific-entity, and specific-relation (see Section 3.3). We also report results without any control condition. Detailed results are shown in Table 1.
Compared to the unconditional setting, the model shows notable improvements in semantic quality under conditional constraints, likely due to the additional guidance that conditions provide. Across all condition types, the model consistently achieves high condition-adherence accuracy, with most values exceeding 80%. Structural constraints generally result in better semantic similarity than semantic-focused conditions, possibly because they help eliminate more irrelevant hypotheses. Among structural conditions, using a fixed format yields the best performance in both semantic and adherence metrics, as it simultaneously constrains entity count, relation count, and logical pattern. While both specific-entity and specific-relation significantly enhance semantic similarity to a similar extent, the model shows a clear preference in condition adherence toward specific-relation.
4.3 Ablation Study
We further analyzed the influence of two proposed components of CtrlHGen, dataset augmentation based on sub-logical decomposition and the reward function.
Sub-logical Decomposition. We evaluate 13 logical patterns on DBpedia-50 using predefined patterns as conditions. The evaluation is conducted under two settings: one with the data augmentation strategy and one without it. As shown in Fig. 5, sub-logical decomposition significantly improves the Jaccard Index, especially for complex patterns involving disjunctions and negations, while maintaining similar Accuracy between two settings. This indicates that the improvement in long logic is due to the enhanced understanding of the internal logical structure rather than relying on external prompts. Notably, improvements also appear on simple patterns (e.g., 1p), indicating the model benefits from decomposing logic into simpler sub-components.
Reward Function. We investigate different reward functions on WN18RR with the "pattern" condition. The results has been shown in Table 2. Reinforcement learning notably improves generalization and reduces accuracy variance compared to supervised learning. Removing Dice and Overlap rewards weakens performance, indicating that Jaccard alone is too strict and may hinder convergence. Excluding the condition-adherence reward slightly improves semantic similarity but harms condition adherence, confirming our reward design effectively balances both objectives.
(a) Jaccard Score
(b) Condition Adherence Accuracy
Figure 5: Results of ablation studies for the sub-logical decomposition
Table 2: Results of ablation studies for the reward function.
Model Semantic Similarity Condition Adherence Average
Jaccard Dice Overlap Accuracy Smatch
CtrlHG(w/o RL) 71.5±0.37 75.8±0.35 83.7±0.33 81.5±0.38 79.0±0.18 78.3
CtrlHG(w/o Dice and Overlap) 74.8±0.34 78.2±0.33 85.1±0.30 90.3±0.25 82.0±0.15 82.1
CtrlHG(w/o Condition Adherence) 77.5±0.33 81.6±0.31 87.8±0.29 68.3±0.46 75.0±0.22 78.0
CtrlHG 77.0±0.34 80.8±0.31 86.8±0.28 93.5±0.24 83.3±0.15 84.3
4.4 Visualization
To evaluate controllability, we sampled 100 hypothesis-observation pairs from the FB15k-237 test set for each category defined by the number of relations (1, 2, or 3) in the reference hypothesis. We compared the number of predicate relations in generated hypotheses under two settings: with and without relation-number constraints. As shown in Fig. 6, without conditional constraints, the model tends to generate hypotheses with a larger number of predicate relations, making it difficult to generate hypotheses with only one relation. However, when conditional constraints are applied, the majority of generated hypotheses align with the expected number of predicates. This experiment further demonstrates the strong controllability of our model.
(a) Without condition constraint
(b) With relation-number condition
Figure 6: Visualization of Relation-number Distribution in Generated Hypotheses
4.5 Case study
To further illustrate controllable hypothesis generation, we present two representative cases from FB15k-237. In the first, the model generates fine-grained hypotheses that cover more relevant artists by controlling the logical structure. In the second, it focuses on specific entities, uncovering both subtle and strong associations between observation and target entities. Even for seemingly unrelated ones (e.g., BAFTA Award), the model uses logical disjunctions to preserve semantic relevance and constraint compliance. The corresponding figure and detailed analysis appear in Appendix C.
5 Conclusion
In summary, this paper introduces a new task of controllable abductive reasoning in knowledge graphs to address the limitation of controllability in the existing method. To tackle the challenges when control generating long and complex logical hypotheses, we propose a data augmentation strategy based on sub-logic decomposition, along with smoother semantic and constraint-adherence reward functions. Experimental results demonstrate that our approach significantly improves the controllability and overall quality of the generated hypotheses.
Appendix A Details for Observation-Hypothesis Sample
Given a knowledge graph $G$ and a predefined logical pattern $P$, the algorithm begins by sampling a random node $v$ and recursively constructs a hypothesis such that $v$ is one of its conclusions and the hypothesis conforms to the logical type specified by $P$. During the recursive process, the algorithm examines the current operation in the hypothesis structure. If the operation is projection, the algorithm randomly selects an incoming edge $(u,r,v)$ of node $v$, then recursively generates a sub-hypothesis rooted at node $u$ according to the corresponding subtype of $P$. If the operation is intersection, the algorithm recursively constructs sub-hypotheses using the same node $v$ for each subtype, since all sub-hypotheses must conclude with $v$. If the operation is union, it applies the recursion to one subtype using node $v$, and to the remaining subtypes using randomly selected nodes. This is because, under union, only one of the sub-hypotheses needs to have $v$ as its conclusion.
For the sub-logic decomposition, we decompose a hypothesis into its sub-logical hypothesis $H_{sub}$ based on the type of reference hypothesis $H$. For example, a logical pattern "inp" can be decomposed into two sublogical patterns "2p". Then we calculate the corresponding conclusions of these two "2p" logical hypotheses respectively as sub-observations, thereby constructing the sub-logic observation-hypothesis set.
Appendix B More experiment Details
For all experiments, we set the learning rate to 1e-5 and use a batch size of 256 during supervised training. The supervised training process consists of two stages. In the first stage, the model is trained for 400 epochs, including a 50-epoch warm-up phase. In the second stage, which involves conditional supervised training, we train for 50 epochs with a 5-epoch warm-up. For reinforcement learning, a smaller batch size of 32 is used, and each group samples 4 candidate answers. The hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 1.0, 0.5, and 0.5, respectively. And then we set $\alpha=0.5$.
In this work, we constrain both observations and hypotheses to be composed solely of entities, relations from the knowledge graph, and logical connectives. To prevent semantic leakage from pre-trained language models, we represent entities and relations using their corresponding IDs in the knowledge graph. As a result, the task is essentially a sequence generation or graph generation problem, rather than a traditional natural language processing task. Accordingly, we adopt the autoregressive GPT-2 architecture as our model, and do not employ more advanced GPT variants, even though they are permissible.
Appendix C More results
C.1 Case study
To further illustrate controllable hypothesis generation, we present two representative cases from FB15k-237. In the first case shown in Fig. 7, the observation is four different types of music { Blues, Jazz, Rhythm_and_blues, Bebop}. By incorporating more complex logical patterns as conditions, the model is able to generate increasingly fine-grained answers. For example, under the simplest "1p" logical pattern, the model can identify their common parent music genre. As the complexity of the logical pattern increases, the model becomes capable of answering more detailed queries, such as identifying artists associated with those music genres. In the second case shown in Fig. 8, it focuses on specific entities. For strongly related entities such as Yahoo, the model is able to identify clear connections with the observation set. Even for entities with weaker relationships, such as two movies, the model can still capture hidden associations between them. Surpringsingly, for seemingly unrelated entities like BAFTA_Award_for_Best_Sound, the model is able to generate high-quality hypotheses by leveraging a logical structure involving the "or" operator, while still adhering to the given constraints.
Observation: Blues, Jazz, Rhythm_and_blues, Bebop Condition1:  Logical Pattern 1p Hypothesis 1:  $H=V_{?}:Parent\_genre(Hard\_bop,V_{?})$ Interpretations 1:  The music genre that originates from the Hard_bop genre. Conclusion 1:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0   Condition2:  Logical Pattern 2p
Hypothesis 2:  $H=V_{?}:Parent\_genre(P_{?},V_{?})\land genre(McCoy\_Tyner,P_{?})$
Interpretations 2:  The musical genre that originates from the genre which is associated with the artist McCoy_Tyner.
Conclusion 2:  Blues, Jazz, Rhythm_and_blues, Bebop.
Jaccard Score:  1.0
  Condition3:  Logical Pattern ip Hypothesis 3:  $H=V_{?}:Parent\_genre(Hard\_bop,V_{?})\land genre(Roy\_Haynes,P_{?})\land genre(McCoy\_Tyner,P_{?})$ Interpretations 3:  The musical genre that originates from the Hard_bop genre and is associated with the artist Roy_Haynes and McCoy_Tyner. Conclusion 3:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0
Figure 7: Case study of Logic Control.
Observation: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay Condition 1:  Entity contains Smokin_Aces Hypothesis 1:  $H=V_{?}:Employ(V_{?},P_{?})\land Film\_crew\_role(Smokin\_Aces,P_{?})$ Interpretations 1:  The organization/company that employs people who have served as a crew member in the film Smokin_Aces. Conclusion 1:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0   Condition 2:  Entity contains The_Phantom
Hypothesis 2:  $H=V_{?}:Employ(V_{?},P_{?})\land Film\_crew\_role(The\_Phantom,P_{?})$
Interpretations 2:  The organization/company that employs people who have served as a crew member in the film The_Phantom.
Conclusion 2:  Blues, Jazz, Rhythm_and_blues, Bebop.
Jaccard Score:  1.0
  Condition 3:  Entity contains Yahoo Hypothesis 3:  $H=V_{?}:Employ(V_{?},P_{?})\land\neg Employed(P_{?},Yahoo)\land Employed(P_{?},Avid\_Technology)$ Interpretations 3:  The organization/company that employs people who have never been employed by Yahoo, but have been employed by Avid_Technology. Conclusion 3:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0   Condition 4:  Entity contains BAFTA_Award_for_Best_Sound Hypothesis 4:  $H=V_{?}:Employ(V_{?},P_{?})\land Nominated_{f}or(P_{?},BAFTA\_Award\_for\_Best\_Sound)\lor Employed(P_{?},Los\_Angeles\_Kings)$ Interpretations 4:  The organization/company that employs people who have been for BAFTA_Award_for_Best_Sound or have been employed by Los_Angeles_Kings. Conclusion 4:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0
Figure 8: Case Study of Entity Semantic Control.
Appendix D Limitation
Our method requires retraining on each individual knowledge graph. As a result, a separate model must be trained for each specific domain, leading to relatively high transfer costs. Additionally, we only explore the model’s performance under single-condition control and do not conduct experiments with composite conditions. However, we believe that our framework is theoretically capable of solving such tasks.
Appendix E Broader Impact
This paper aims to promote the development of abductive reasoning in knowledge graphs. Our work has contributed to the field of deep learning and has brought about many potential social consequences. It may play an important role in fields such as clinical diagnosis or scientific discovery. However, we believe there are no negative impacts that need clarification.